OpenStudy (anonymous):

Find the equation of the tangent to the graph at the indicated point. f(x) = x2 − 3; a = 9 Y=

OpenStudy (anonymous):

[f(x) = x^2 − 3 f'(x) =lim [(x+h)^2-3 - (x^2 − 3)]/h h=>0 f'(x) =lim [(x^2+2xh+h^2-3) - (x^2 − 3)]/h h=>0 f'(x) =lim [(x^2+2xh+h^2-3) - x^2 + 3)]/h h=>0 f'(x) =lim [(2xh+h^2)]/h h=>0 f'(x) =lim [(2x+h)] h=>0 f'(x) = [(2x+0)] f'(x) = 2x now at point a=9 f'(9) = 2(9)=18 .....ans

OpenStudy (anonymous):

ok thnx,n gud luck

OpenStudy (anonymous):

No thanks you!!